OpenStudy (ashking1):

Which is the exact value of the expression sqrt96+sqrt150-sqrt54

sam (.sam.):

\[\sqrt{96}+\sqrt{150}-\sqrt{54}\] \[\sqrt{16}\sqrt{6}+\sqrt{25}\sqrt{6}-\sqrt{9}\sqrt{6}\] \[4\sqrt{6}+5\sqrt{6}-3\sqrt{6}\] Add them up

OpenStudy (ashking1):

6sqrt6

sam (.sam.):

yes

OpenStudy (ashking1):

ok thanks